OpenStudy (anonymous):

a scientist prepares a solution by mixing 25ml of solution g with 50ml of solution h resulting in a solution with a mass of 125 grams . he then combined 45ml of solution g with 20 ml of solution h which procedures a solution with a mass of 120 grams in the equation below g represents the concentration of solution g and h represents the concentration of solution h . what is the concentration of solution h , in grams/ml

OpenStudy (anonymous):

the equation is like this 25g+50h=125 and 45g+20h=120

OpenStudy (wolfe8):

Something seems weird to me about this question. It seems that g and h are not concentrations but density. Let me think it over

OpenStudy (anonymous):

ok

OpenStudy (wolfe8):

But I guess if it says that, then solving the simultaneous equation should give you the concentrations for g and h.

OpenStudy (anonymous):

25g+50h=125 and 45g+20h=120

OpenStudy (wolfe8):

Ah it makes sense now. This is a math question not really chemistry. I took it as being complicated. Just solve the simultaneous equation problem for h.

OpenStudy (anonymous):

so i need to ask the math people

OpenStudy (wolfe8):

Yes but since it's already here I'll help you. Are you familiar with the methods of solving them? You have substitution and elimination. http://www.allaboutcircuits.com/vol_5/chpt_4/11.html

OpenStudy (anonymous):

yes i know how to do them

OpenStudy (wolfe8):

Right. So is there anything else you need help with?

OpenStudy (anonymous):

i dont know let me try again 2 do it